OpenStudy (anonymous):

calculate: lim as x approaches pi/4 = [(x - pi/4)^2 / (tan x - 1)^2] .

hartnn (hartnn):

the first thing that should pop up in the mind is to substitute \(u = x-\pi/4 \\ as, x\to \pi/4, u\to 0 \) make this substitution! and tell me what u get ?

OpenStudy (anonymous):

lim as u approaches 0 = [(u^2) / (tan x-1)^2] ?

hartnn (hartnn):

you have to change tan x as well! :) as, u = x- pi/4 x = u + pi/4 right ? so tan x changes to tan (u+ pi/4) got it ?

OpenStudy (anonymous):

ya :D. then?

hartnn (hartnn):

expand tan (u+pi/4) do you know the formula for tan (A+B) ?

hartnn (hartnn):

also put tan pi/4 = 1 :)

OpenStudy (anonymous):

it will be : [(tan x +tan pi/4)/ (1-tan x.tan pi/4)] ?

OpenStudy (anonymous):

sorry, imean it will be : [(tan u +tan pi/4)/ (1-tan u.tan pi/4)] ?

hartnn (hartnn):

yes

hartnn (hartnn):

now put tan pi/4 = 1

OpenStudy (anonymous):

(tan u + 1)/(1-tan u) ?

hartnn (hartnn):

yes now we had tan x -1 in the denominator so can you try to simplify (tan u + 1)/(1-tan u) -1 = ... ?

OpenStudy (anonymous):

will it be (2 tan u)/(1-tan u) ?

hartnn (hartnn):

correct! plug this in your function

OpenStudy (anonymous):

(u^2)/(2tan u/1-tan u)^2 ..

hartnn (hartnn):